# Tracing Quenching in Nearby Galaxies Through Inner Surface Mass Density and Cold Gas Content

### v1

@@ 2.5 Overlap Sample
Also shown in Figure 2 are the distributions for the overlap sample. In general, the overlap sample has similar distributions to our HI and H2 samples. One notable discrepancy is in stellar mass. This is made clearer in Figure 4, which compares the distributions for the HI, H2, and overlap samples directly. More than 75% of galaxies in our full H2 sample are also in the HI sample, but we find that our full H2 sample is slightly biased toward lower masses compared to both the HI and overlap samples. This can be attributed to the differing distributions in $M_{\ast}$ between the xGASS and xCOLDGASS samples (Figure 2). We also note that galaxies with masses between $10<\log(M_{\ast}/M_{\odot})<10.25$ appear to be underrepresented in our HI sample compared to the xGASS sample (Figure 2). These discrepancies notwithstanding, we show below that they do not affect our results.
@@
3 Results
3.1 Statistical Methods
Since our samples include objects with low SFR (and presumably low gas content), it is important that we properly treat objects in our samples that have non-detections in HI and/or H2. It is common to use statistical methods such as survival analysis or spectral stacking to robustly treat upper limits for the purposes of including them in data analysis. For a recent overview of such methods in astronomy, see Saintonge & Catinella (2022). Within bins of $\Sigma_{1}$ or $\Delta\Sigma_{1}$, the detection rate was as low as 20%. The median of the distribution is not meaningful in such cases. The Kaplan-Meier (KM) estimator, a non-parametric statistic used in survival analysis, can provide feasible sample statistics down to a detection rate of approximately 20%, albeit with greater uncertainty. In order to include upper limits in our analysis, we elect to use the KM estimator to compute means in the gas fraction within bins of the independent variable in our analyses, adhering to the methods of Feigelson & Nelson (1985). We simulated uncertainty in the gas fraction by bootstrapping; the data was resampled 1000 times with replacement, and the survival function and resulting binned average gas fractions were recomputed for each trial. We use the 5th and 95th percentiles in the distribution of gas fraction means for the 1000 trials to assign a measure of confidence to the “true” gas fraction means estimated via the KM estimator.
Quantifying the slope or scatter of a linear trend is also challenging in the presence of upper limits, when traditional methods are no longer valid. Consequently, we apply the Bayesian hierarchical model for linear regression implemented in the IDL procedure LINMIX_ERR (Kelly 2007). This method is capable of fitting a straight line to data which includes non-detections in the $y$ variable, making it a suitable choice for our study. In addition to linear regression parameters, LINMIX_ERR returns the Gaussian intrinsic scatter of the dependent variable about the regression line. This method has been used widely in astronomy (e.g. Bentz et al. 2013; McConnell & Ma 2013), and it has recently been applied to xCOLDGASS data (Hagedorn et al. 2024). For details about the method itself, see Kelly (2007).
@@ 3.1 Statistical Methods
Figure 6: The dependence of (a) $f_{\mathrm{HI}}$ and (b) $f_{\mathrm{H2}}$ on the relative parameter $\Delta\Sigma_{1}$, coded by specific SFR. The black points represent the mean $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$ values computed using survival analysis techniques within bins of $\Delta\Sigma_{1}\approx 0.12$ dex. The red points represent the mean values within the overlap sample. Error bars are the 5th and 95th percentiles in the bootstrap distribution as described at the beginning of Section 3. (c-d) For both samples, we plot the fraction of galaxies with gas detections and the fraction without as a function of $\Delta\Sigma_{1}$, relative to the total number of galaxies within a given bin of $\Delta\Sigma_{1}$. For (c) the HI sample, we count marginal detections as detected. (e-f) The fraction of quenched and star-forming galaxies as a function of $\Delta\Sigma_{1}$. Galaxies are defined as star-forming or quenched according to the SFR-$M_{\ast}$ line defined by Woo et al. 2015. In plots (c-f), error estimates are obtained via bootstrapping. The thinner, transparent lines indicate the relevant fractions in the overlap sample. The trends shown in (c-f) for the full samples are identical (within the uncertainties) to those found using the overlap sample.
@@
3.2 $\Sigma_{1}$ and Cold Gas Fraction
In the top row of Figure 5, we plot $f_{\mathrm{HI}}$ vs. stellar mass and structural parameters $\mu_{\ast}$ and $\Sigma_{1}$. We confirm previously established findings that $f_{\mathrm{HI}}$ is negatively correlated with $M_{\ast}$ and $\mu_{\ast}$ for nearby galaxies (e.g. Catinella et al. 2010; Huang et al. 2012; Kauffmann et al. 2012; Brown et al. 2015; Catinella et al. 2018; Hunt et al. 2020). We find that there is a similarly strong dependence of $f_{\mathrm{HI}}$ on $\Sigma_{1}$.
In agreement with Saintonge et al. (2017), we observe negative correlations between $f_{\mathrm{H2}}$ and both $M_{\ast}$ and $\mu_{\ast}$ in the bottom row of Figure 5. We similarly note considerable scatter in the dependence of $f_{\mathrm{H2}}$ on $M_{\ast}$, with an intrinsic scatter of approximately $0.43$ dex.
@@ 3.2 $\Sigma_{1}$ and Cold Gas Fraction
We find that $\log f_{\mathrm{HI}}$ also decreases linearly with $\log\Sigma_{1}$ for $\Sigma_{1}>10^{9}\ M_{\odot}\ \mathrm{kpc}^{-2}$, with a slope of $1.11\pm 0.07$ and an intrinsic scatter of $0.42\pm 0.01\ \mathrm{dex}$. Interestingly, $\log f_{\mathrm{HI}}$ vs. $\log\Sigma_{1}$ does not exhibit nonlinear behavior to the extent we observe for $\log f_{\mathrm{H2}}$ vs. $\log\Sigma_{1}$. We elaborate on this observation in Section 4.1.
3.3 $\Delta\Sigma_{1}$ and Cold Gas Fraction
In Figure 6, we show the distribution of our samples in the plane of gas fraction vs. $\Delta\Sigma_{1}$ color-coded by sSFR. The parameter $\Delta\Sigma_{1}$ is defined as the offset in $\log\Sigma_{1}$ relative to the “structural valley”, a feature in the plot of $\log\Sigma_{1}$ vs. $\log M_{\ast}$ which generally separates bulge-dominated from disk-dominated galaxies at fixed mass (Luo et al. 2020). The bin widths are approximately $0.12$ dex in $\Delta\Sigma_{1}$ in both samples.
We observe an ‘elbow’ pattern in the plane of $f_{\mathrm{H2}}$ vs. $\Delta\Sigma_{1}$, a shape that often emerges in the planes of star formation or color vs. $\Sigma_{1}$ or $\Delta\Sigma_{1}$ (Fang et al. 2013; Barro et al. 2017; Lee et al. 2018; Luo et al. 2020; Guo et al. 2021). Following the trend in KM means, the H2 fraction exhibits minor variation until just before $\Delta\Sigma_{1}=0$, at which point the mean fraction decreases by an order of magnitude over an interval of $\Delta\Sigma_{1}\approx 0.4$ dex. This trend is also seen in the overlap sample (Figure 6).
@@ 3.3 $\Delta\Sigma_{1}$ and Cold Gas Fraction
In agreement with Luo et al. (2020), we find that there is a critical point marked by $\Delta\Sigma_{1}=0$ after which point star formation is generally suppressed. This is illustrated in Figure 6, where the shading of points indicates that sSFR declines with increasing $\Delta\Sigma_{1}$ and with decreasing $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$. Galaxies with lower sSFR are found preferentially where $\Delta\Sigma_{1}>0$.
To quantify this, we employ the star-forming criterion of Woo et al. (2015) to define the fraction of quenched objects as a function of $\Delta\Sigma_{1}$ (Figure 6). Among galaxies with $\Delta\Sigma_{1}>0$, we find that $65.6\%$ of our HI sample and $62.3\%$ of our H2 sample are quiescent. In contrast, only $7.2\%$ and $5.6\%$ of galaxies with $\Delta\Sigma_{1}<0$ are quiescent in our HI and H2 samples, respectively. Consequently, we contend that $\Delta\Sigma_{1}=0$ approximately marks a transition from star-forming to quiescent galaxies.
3.4 AGN Activity
@@
While AGN are overrepresented in our sample compared to the Luo et al. (2020) sample (Section 2.4), the general trend is the same between their sample and ours; there is an increase in AGN activity for $\Delta\Sigma_{1}>0$, and little AGN activity for $\Delta\Sigma_{1}<0$. Within the Luo et al. (2020) sample of galaxies – after we impose the constraints that $M_{\ast}>10^{10}M_{\odot}$ and $0.02\leq z\leq 0.05$ to be consistent with our samples – approximately $50\%$ of the galaxies located above the $\Sigma_{1}-M_{\ast}$ structural valley ($\Delta\Sigma_{1}>0$) are AGN candidates, compared to $\approx 12\%$ of galaxies below. This suggests that AGN candidates are possibly overrepresented in our samples for $\Delta\Sigma_{1}>0$. Regardless, in both our sample and the parent sample, we can conclude that the threshold $\Delta\Sigma_{1}>0$ is associated with a greater incidence of AGN candidates relative to the region $\Delta\Sigma_{1}<0$.
@@
4 Discussion
4.1 The Relationship between $\Sigma_{1}$ and Cold Gas
@@
As the detected sample consists predominantly of main-sequence galaxies, this observation could be explained by appealing to the typical galactic distribution of cold gas in massive disk galaxies. HI is accreted into the outskirts of galaxies via gas infall from the intergalactic medium or cosmic filaments, after which point it can replenish the inner parts of the galaxy and fuel star formation via conversion into the molecular phase (see e.g. Sancisi et al. (2008) for a review of cold gas accretion). As a result, the centers of nearby spiral galaxies are dominated by H2 gas, while the outskirts are dominated by HI gas (Young & Scoville 1991; Bigiel et al. 2008; Leroy et al. 2008). Since $\Sigma_{1}$ probes the central regions of galaxies, we expect $f_{\mathrm{H2}}$ to exhibit greater sensitivity to $\Sigma_{1}$ than $f_{\mathrm{HI}}$. This is further supported by the more direct connection between $\Delta\Sigma_{1}$ and $f_{\mathrm{H2}}$ compared to $f_{\mathrm{HI}}$, which is discussed in depth in the next subsection. However, it should be noted that the mild ‘elbow’ shape observed in the plane of $f_{\mathrm{H2}}$ vs. $\Sigma_{1}$ could be a byproduct of the tighter correlation between $f_{\mathrm{H2}}$ and SFR, and of the nonlinear behavior which has been reported between $\Sigma_{1}$ and sSFR (e.g. Barro et al. 2013).
4.2 $\Delta\Sigma_{1}$ as an Indictor of Evolution Towards Quenching
@@
Previous literature has proposed that galaxies evolve towards higher $\Sigma_{1}$ and that galaxies quench at some mass-dependent threshold (Fang et al. 2013; van Dokkum et al. 2015; Barro et al. 2017). Luo et al. (2020) contended that $\Delta\Sigma_{1}=0$ is the demarcation between pseudo-bulges, which lie to the left of $\Delta\Sigma_{1}=0$, and classical bulges, which lie to the right. They observed a “bimodality” in the distribution of $\Delta\Sigma_{1}$ vs. global sSFR, whereby galaxies to the left of $\Delta\Sigma_{1}=0$ have systematically greater values of sSFR, while galaxies to the right have lower values, but with some range in sSFR. Luo et al. (2020) raised the possibility that some galaxies evolve from pseudo- to classical-type bulges, or from low to high $\Delta\Sigma_{1}$, and as such they offered an explanation for the greater range of sSFR values to the right of $\Delta\Sigma_{1}=0$ as the result of differences in structure and stellar populations during evolution towards quenching.
@@
We observe a decrease in both $f_{\mathrm{H2}}$ and $f_{\mathrm{HI}}$ across $\Delta\Sigma_{1}=0$, and a corresponding decrease in sSFR (Figure 6). While $f_{\mathrm{HI}}$ exhibits an approximately linear decline with $\Delta\Sigma_{1}$, $f_{\mathrm{H2}}$ decreases nonlinearly with $\Delta\Sigma_{1}$, particularly across the $\Delta\Sigma_{1}$ zero point as discussed in Section 3.3. This makes sense in light of the fact that the central gas reservoir is dominated by molecular gas, as discussed in Section 4.1. Additionally, there is evidence to suggest that H2 content is more tightly correlated with galaxy quiescence than HI. Zhang et al. (2019) observed evidence for significant HI gas reservoirs in nearby, massive quiescent galaxies which were classified as centrals with disk morphology. A significant proportion of their sample had HI detections but were generally depleted in H2 gas.
@@
In accordance with Luo et al. (2020), we find that $\Delta\Sigma_{1}$, like $\Sigma_{1}$, is indicative of quenching, and furthermore, that $\Delta\Sigma_{1}$ may be indicative of evolution towards quenching. It is tempting to consider that the mean trend of $\Delta\Sigma_{1}$ vs. $f_{\mathrm{H2}}$ in Figures 6 and 7 constitutes an evolutionary track; in the remainder of this subsection, we show that this is plausible.
@@
Ma et al. (2022) employed cosmological simulations to examine how $\Delta\Sigma_{1}$ correlates with offsets in cold gas mass—specifically, $\Delta\log M_{\mathrm{H{\small I}MS}}$ and $\Delta\log M_{\mathrm{H_{2}MS}}$—where each is defined as the deviation in either $M_{\mathrm{HI}}$ or $M_{\mathrm{H_{2}}}$ from the corresponding $_{\mathrm{gas}}MM_{\ast}$ “main sequence.” In their work, this main sequence is constructed as the median value of $\log M_{\mathrm{gas}}$ in bins of $\log M_{\ast}$ for star-forming galaxies (SFGs); similarly, $\Delta\Sigma_{1}$ is defined as the deviation from the median $\Sigma_{1}$ in bins of $M_{\ast}$ for SFGs. Our results reproduce the general trends of their Figure 11, although we find a steeper correlation between H2 content and $\Delta\Sigma_{1}$ than reported by Ma et al. (2022). While $f_{\mathrm{H2}}$ declines by $\sim 1$ dex between $\Delta\Sigma_{1}\approx-0.2$ and $0.2$ dex in our sample, Ma et al. (2022) find that $\log M_{\mathrm{H_{2}MS}}$ decreases by only $\sim 0.25$ dex in this same range in $\Delta\Sigma_{1}$. We also observe a flatter trend between HI gas fraction and $\Delta\Sigma_{1}$ than they report. We attribute these discrepancies partly to our different definition of $\Delta\Sigma_{1}$ and, to our choice of using cold gas fractions (i.e., $M_{\mathrm{H_{2}}}/M_{\ast}$) rather than a cold gas main sequence. Our primary goal is to assess $\Delta\Sigma_{1}$ as an indicator of quenching, so employing the definition from Luo et al. (2020) – which is independent of any star-forming or quenched designation – offers a potentially less biased parameterization for tracking the depletion of cold gas during quenching. Regardless of the technical differences, our work provides observational evidence supporting the conclusion of Ma et al. (2022).
@@
However, we cannot rule out the possibility that the apparent differences in how $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$ depend on $\Delta\Sigma_{1}$ are merely byproducts of other correlations – namely the bimodal relationship between $\Delta\Sigma_{1}$ and sSFR, and the different dependencies of molecular and atomic gas fractions on sSFR. In particular, $f_{\mathrm{H2}}$ correlates more strongly with sSFR than does $f_{\mathrm{HI}}$ (e.g. Saintonge & Catinella 2022). Moreover, because H2 is typically more concentrated in galaxy centers, the differing trends observed between $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$ with $\Delta\Sigma_{1}$ could simply reflect quenching mechanisms associated with central bulge growth and with the resulting reduction in global sSFR, rather than physical differences in HI vs. H2 gas depletion. In any case, our results support that the interplay of these three parameters is important for the quenching of massive galaxies at fixed mass.
4.3 The Connection to AGN Activity

### v2

@@ II.5 Overlap Sample
Also shown in Figure 2 are the distributions for the overlap sample. In general, the overlap sample has similar distributions to our HI and H2 samples. One notable discrepancy is in stellar mass. This is made clearer in Figure 4, which compares the distributions for the HI, H2, and overlap samples directly. More than 75% of galaxies in our full H2 sample are also in the HI sample, but we find that our full H2 sample is slightly biased toward lower masses compared to both the HI and overlap samples. This can be attributed to the differing distributions in $M_{\ast}$ between the xGASS and xCOLDGASS samples (Figure 2). We also note that galaxies with masses between $10<\log(M_{\ast}/M_{\odot})<10.25$ appear to be underrepresented in our HI sample compared to the xGASS sample (Figure 2).
To further quantify any differences between our HI and H2 samples, we use the two-sample Kolmogorov-Smirnov (K-S) statistic to estimate similarity in the distributions of $M_{\ast}$, $\mu_{\ast}$, $\Sigma_{1}$, and SFR between our two samples. Because the significant overlap between our samples violates the assumption of independence in the naïve application of the K-S test, we perform a permutation test on the K-S statistic that preserves the shared data and randomly reassigns the non-shared data (Praestgaard1995). We calculate the K-S statistic ($D$) and its empirical $p$ value after $10,000$ permutations. We choose a significance level of $\alpha=0.05$ for testing the null hypothesis that the HI and H2 distributions are identical for the property in consideration.
We find no statistically significant differences between the HI and H2 distributions of SFR ($p=0.10$) and $\mu_{\ast}$ ($p=0.12$). We find statistically significant, but small differences between the HI and H2 distributions for $\Sigma_{1}$ ($D=0.10$, $p=0.001$) and $M_{\ast}$ ($D=0.10$, $p=0.003$). These differences notwithstanding, we show below that they do not affect our results.
@@
III Results
III.1 Statistical Methods
Since our samples include objects with low SFR (and presumably low gas content), it is important that we properly treat objects in our samples that have non-detections in HI and/or H2. It is common to use statistical methods such as survival analysis or spectral stacking to robustly treat upper limits for the purposes of including them in data analysis. For a recent overview of such methods in astronomy, see Saintonge22. Within bins of $\Sigma_{1}$ or $\Delta\Sigma_{1}$, the detection rate was as low as 20%. The median of the distribution is not meaningful in such cases. The Kaplan-Meier (KM) estimator, a non-parametric statistic used in survival analysis, can provide feasible sample statistics down to a detection rate of approximately 20%, albeit with greater uncertainty. In order to include upper limits in our analysis, we elect to use the KM estimator to compute means in the gas fraction within bins of the independent variable in our analyses, adhering to the methods of Feigelson85. We simulated uncertainty in the gas fraction by bootstrapping; the data was resampled 1000 times with replacement, and the survival function and resulting binned average gas fractions were recomputed for each trial. We use the 5th and 95th percentiles in the distribution of gas fraction means for the 1000 trials to assign a measure of confidence to the “true” gas fraction means estimated via the KM estimator.
Quantifying the slope or scatter of a linear trend is also challenging in the presence of upper limits, when traditional methods are no longer valid. Consequently, we apply the Bayesian hierarchical model for linear regression implemented in the IDL procedure LINMIX_ERR (LINMIX2007). This method is capable of fitting a straight line to data which includes non-detections in the $y$ variable, making it a suitable choice for our study. In addition to linear regression parameters, LINMIX_ERR returns the Gaussian intrinsic scatter of the dependent variable about the regression line. This method has been used widely in astronomy (e.g. Bentz_2013; McConnell_Ma_2013), and it has recently been applied to xCOLDGASS data (Hagedorn2024). For details about the method itself, see LINMIX2007.
@@ III.1 Statistical Methods
Figure 6: The dependence of (a) $f_{\mathrm{HI}}$ and (b) $f_{\mathrm{H2}}$ on the relative parameter $\Delta\Sigma_{1}$, coded by specific SFR. The black points represent the mean $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$ values computed using survival analysis techniques within bins of $\Delta\Sigma_{1}\approx 0.12$ dex. The red points represent the mean values within the overlap sample. Error bars are the 5th and 95th percentiles in the bootstrap distribution as described at the beginning of Section III. (c-d) For both samples, we plot the fraction of galaxies with gas detections and the fraction without as a function of $\Delta\Sigma_{1}$, relative to the total number of galaxies within a given bin of $\Delta\Sigma_{1}$. For (c) the HI sample, we count marginal detections as detected. (e-f) The fraction of quenched and star-forming galaxies as a function of $\Delta\Sigma_{1}$. Galaxies are defined as star-forming or quenched according to the SFR-$M_{\ast}$ line defined by Woo15. In plots (c-f), error estimates are obtained via bootstrapping. The thinner, transparent lines indicate the relevant fractions in the overlap sample. The trends shown in (c-f) for the full samples are identical (within the uncertainties) to those found using the overlap sample.
@@
III.2 $\Sigma_{1}$ and Cold Gas Fraction
In the top row of Figure 5, we plot $f_{\mathrm{HI}}$ vs. stellar mass and structural parameters $\mu_{\ast}$ and $\Sigma_{1}$. We confirm previously established findings that $f_{\mathrm{HI}}$ is negatively correlated with $M_{\ast}$ and $\mu_{\ast}$ for nearby galaxies (e.g. Catinella10; Huang12; Kauffmann12; Brown15; Catinella18; Hunt20). We find that there is a similarly strong dependence of $f_{\mathrm{HI}}$ on $\Sigma_{1}$.
In agreement with Saintonge17, we observe negative correlations between $f_{\mathrm{H2}}$ and both $M_{\ast}$ and $\mu_{\ast}$ in the bottom row of Figure 5. We similarly note considerable scatter in the dependence of $f_{\mathrm{H2}}$ on $M_{\ast}$, with an intrinsic scatter of approximately $0.43$ dex.
@@ III.2 $\Sigma_{1}$ and Cold Gas Fraction
We find that $\log f_{\mathrm{HI}}$ also decreases linearly with $\log\Sigma_{1}$ for $\Sigma_{1}>10^{9}\ M_{\odot}\ \mathrm{kpc}^{-2}$, with a slope of $1.11\pm 0.07$ and an intrinsic scatter of $0.42\pm 0.01\ \mathrm{dex}$. Interestingly, $\log f_{\mathrm{HI}}$ vs. $\log\Sigma_{1}$ does not exhibit nonlinear behavior to the extent we observe for $\log f_{\mathrm{H2}}$ vs. $\log\Sigma_{1}$. We elaborate on this observation in Section IV.1.
III.3 $\Delta\Sigma_{1}$ and Cold Gas Fraction
In Figure 6, we show the distribution of our samples in the plane of gas fraction vs. $\Delta\Sigma_{1}$ color-coded by sSFR. The parameter $\Delta\Sigma_{1}$ is defined as the offset in $\log\Sigma_{1}$ relative to the “structural valley”, a feature in the plot of $\log\Sigma_{1}$ vs. $\log M_{\ast}$ which generally separates bulge-dominated from disk-dominated galaxies at fixed mass (Luo20). The bin widths are approximately $0.12$ dex in $\Delta\Sigma_{1}$ in both samples.
We observe an ‘elbow’ pattern in the plane of $f_{\mathrm{H2}}$ vs. $\Delta\Sigma_{1}$, a shape that often emerges in the planes of star formation or color vs. $\Sigma_{1}$ or $\Delta\Sigma_{1}$ (Fang13; Barro17; Lee2018; Luo20; Guo2021). Following the trend in KM means, the H2 fraction exhibits minor variation until just before $\Delta\Sigma_{1}=0$, at which point the mean fraction decreases by an order of magnitude over an interval of $\Delta\Sigma_{1}\approx 0.4$ dex. This trend is also seen in the overlap sample (Figure 6).
@@ III.3 $\Delta\Sigma_{1}$ and Cold Gas Fraction
In agreement with Luo20, we find that there is a critical point marked by $\Delta\Sigma_{1}=0$ after which point star formation is generally suppressed. This is illustrated in Figure 6, where the shading of points indicates that sSFR declines with increasing $\Delta\Sigma_{1}$ and with decreasing $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$. Galaxies with lower sSFR are found preferentially where $\Delta\Sigma_{1}>0$.
To quantify this, we employ the star-forming criterion of Woo15 to define the fraction of quenched objects as a function of $\Delta\Sigma_{1}$ (Figure 6). Among galaxies with $\Delta\Sigma_{1}>0$, we find that $65.6\%$ of our HI sample and $62.3\%$ of our H2 sample are quiescent. In contrast, only $7.2\%$ and $5.6\%$ of galaxies with $\Delta\Sigma_{1}<0$ are quiescent in our HI and H2 samples, respectively. Consequently, we contend that $\Delta\Sigma_{1}=0$ approximately marks a transition from star-forming to quiescent galaxies.
III.4 AGN Activity
@@
While AGN are overrepresented in our sample compared to the Luo20 sample (Section II.4), the general trend is the same between their sample and ours; there is an increase in AGN activity for $\Delta\Sigma_{1}>0$, and little AGN activity for $\Delta\Sigma_{1}<0$. Within the Luo20 sample of galaxies – after we impose the constraints that $M_{\ast}>10^{10}M_{\odot}$ and $0.02\leq z\leq 0.05$ to be consistent with our samples – approximately $50\%$ of the galaxies located above the $\Sigma_{1}-M_{\ast}$ structural valley ($\Delta\Sigma_{1}>0$) are AGN candidates, compared to $\approx 12\%$ of galaxies below. This suggests that AGN candidates are possibly overrepresented in our samples for $\Delta\Sigma_{1}>0$. Regardless, in both our sample and the parent sample, we can conclude that the threshold $\Delta\Sigma_{1}>0$ is associated with a greater incidence of AGN candidates relative to the region $\Delta\Sigma_{1}<0$.
@@
IV Discussion
IV.1 The Relationship between $\Sigma_{1}$ and Cold Gas
@@
As the detected sample consists predominantly of main-sequence galaxies, this observation could be explained by appealing to the typical galactic distribution of cold gas in massive disk galaxies. HI is accreted into the outskirts of galaxies via gas infall from the intergalactic medium or cosmic filaments, after which point it can replenish the inner parts of the galaxy and fuel star formation via conversion into the molecular phase (see e.g. Sancisi2008 for a review of cold gas accretion). As a result, the centers of nearby spiral galaxies are dominated by H2 gas, while the outskirts are dominated by HI gas (Young91; Bigiel2008; Leroy2008). Since $\Sigma_{1}$ probes the central regions of galaxies, we expect $f_{\mathrm{H2}}$ to exhibit greater sensitivity to $\Sigma_{1}$ than $f_{\mathrm{HI}}$. This is further supported by the more direct connection between $\Delta\Sigma_{1}$ and $f_{\mathrm{H2}}$ compared to $f_{\mathrm{HI}}$, which is discussed in depth in the next subsection. However, it should be noted that the mild ‘elbow’ shape observed in the plane of $f_{\mathrm{H2}}$ vs. $\Sigma_{1}$ could be a byproduct of the tighter correlation between $f_{\mathrm{H2}}$ and SFR, and of the nonlinear behavior which has been reported between $\Sigma_{1}$ and sSFR (e.g. Barro13).
IV.2 $\Delta\Sigma_{1}$ as an Indictor of Evolution Towards Quenching
@@
Previous literature has proposed that galaxies evolve towards higher $\Sigma_{1}$ and that galaxies quench at some mass-dependent threshold (Fang13; vanDokkum2015; Barro17). Luo20 contended that $\Delta\Sigma_{1}=0$ is the demarcation between pseudo-bulges, which lie to the left of $\Delta\Sigma_{1}=0$, and classical bulges, which lie to the right. They observed a “bimodality” in the distribution of $\Delta\Sigma_{1}$ vs. global sSFR, whereby galaxies to the left of $\Delta\Sigma_{1}=0$ have systematically greater values of sSFR, while galaxies to the right have lower values, but with some range in sSFR. Luo20 raised the possibility that some galaxies evolve from pseudo- to classical-type bulges, or from low to high $\Delta\Sigma_{1}$, and as such they offered an explanation for the greater range of sSFR values to the right of $\Delta\Sigma_{1}=0$ as the result of differences in structure and stellar populations during evolution towards quenching.
@@
We observe a decrease in both $f_{\mathrm{H2}}$ and $f_{\mathrm{HI}}$ across $\Delta\Sigma_{1}=0$, and a corresponding decrease in sSFR (Figure 6). While $f_{\mathrm{HI}}$ exhibits an approximately linear decline with $\Delta\Sigma_{1}$, $f_{\mathrm{H2}}$ decreases nonlinearly with $\Delta\Sigma_{1}$, particularly across the $\Delta\Sigma_{1}$ zero point as discussed in Section III.3. This makes sense in light of the fact that the central gas reservoir is dominated by molecular gas, as discussed in Section IV.1. Additionally, there is evidence to suggest that H2 content is more tightly correlated with galaxy quiescence than HI. Zhang2019 observed evidence for significant HI gas reservoirs in nearby, massive quiescent galaxies which were classified as centrals with disk morphology. A significant proportion of their sample had HI detections but were generally depleted in H2 gas.
@@
In accordance with Luo20, we find that $\Delta\Sigma_{1}$, like $\Sigma_{1}$, is indicative of quenching, and furthermore, that $\Delta\Sigma_{1}$ may be indicative of evolution towards quenching. It is tempting to consider that the mean trend of $\Delta\Sigma_{1}$ vs. $f_{\mathrm{H2}}$ in Figures 6 and 7 constitutes an evolutionary track; in the remainder of this subsection, we show that this is plausible.
@@
Ma2022 employed cosmological simulations to examine how $\Delta\Sigma_{1}$ correlates with offsets in cold gas mass—specifically, $\Delta\log M_{\mathrm{H{\small I}MS}}$ and $\Delta\log M_{\mathrm{H_{2}MS}}$—where each is defined as the deviation in either $M_{\mathrm{HI}}$ or $M_{\mathrm{H_{2}}}$ from the corresponding $_{\mathrm{gas}}MM_{\ast}$ “main sequence.” In their work, this main sequence is constructed as the median value of $\log M_{\mathrm{gas}}$ in bins of $\log M_{\ast}$ for star-forming galaxies (SFGs); similarly, $\Delta\Sigma_{1}$ is defined as the deviation from the median $\Sigma_{1}$ in bins of $M_{\ast}$ for SFGs. Our results reproduce the general trends of their Figure 11, although we find a steeper correlation between H2 content and $\Delta\Sigma_{1}$ than reported by Ma2022. While $f_{\mathrm{H2}}$ declines by $\sim 1$ dex between $\Delta\Sigma_{1}\approx-0.2$ and $0.2$ dex in our sample, Ma2022 find that $\log M_{\mathrm{H_{2}MS}}$ decreases by only $\sim 0.25$ dex in this same range in $\Delta\Sigma_{1}$. We also observe a flatter trend between HI gas fraction and $\Delta\Sigma_{1}$ than they report. We attribute these discrepancies partly to our different definition of $\Delta\Sigma_{1}$ and, to our choice of using cold gas fractions (i.e., $M_{\mathrm{H_{2}}}/M_{\ast}$) rather than a cold gas main sequence. Our primary goal is to assess $\Delta\Sigma_{1}$ as an indicator of quenching, so employing the definition from Luo20 – which is independent of any star-forming or quenched designation – offers a potentially less biased parameterization for tracking the depletion of cold gas during quenching. Regardless of the technical differences, our work provides observational evidence supporting the conclusion of Ma2022.
@@
However, we cannot rule out the possibility that the apparent differences in how $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$ depend on $\Delta\Sigma_{1}$ are merely byproducts of other correlations – namely the bimodal relationship between $\Delta\Sigma_{1}$ and sSFR, and the different dependencies of molecular and atomic gas fractions on sSFR. In particular, $f_{\mathrm{H2}}$ correlates more strongly with sSFR than does $f_{\mathrm{HI}}$ (e.g. Saintonge22). Moreover, because H2 is typically more concentrated in galaxy centers, the differing trends observed between $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$ with $\Delta\Sigma_{1}$ could simply reflect quenching mechanisms associated with central bulge growth and with the resulting reduction in global sSFR, rather than physical differences in HI vs. H2 gas depletion. In any case, our results support that the interplay of these three parameters is important for the quenching of massive galaxies at fixed mass.
IV.3 The Connection to AGN Activity